# A Reproducible Framework for Neural Topic Modeling in Focus Group Analysis

### v1

@@ 5.3 Comparison with LDA Baseline
We compared our final hierarchical BERTopic model to a systematically tuned LDA baseline. Grid search across six configurations identified 7 topics as optimal for LDA, achieving coherence of 0.5464, marginally lower than our BERTopic model (9 topics, coherence = 0.558). This small difference suggests comparable performance on semantic coherence.
However, the minimal agreement between topic assignments (NMI = 0.122, ARI = 0.080) indicates fundamentally different thematic structures. This divergence suggests that architectural differences (LDA’s bag-of-words assumptions versus BERTopic’s transformer-based embeddings) lead to qualitatively distinct topic decompositions. For conversational data where meaning depends on discourse context (e.g., "je n’ai pas confiance" referencing different trust domains), contextual embeddings may capture nuances that statistical co-occurrence patterns miss.
The near-independence of topic assignments (NMI = 0.122) suggests these methods may be complementary rather than competitive, capturing different facets of discourse structure. Future work with parallel human evaluation of both LDA and BERTopic topics could clarify whether these differences reflect genuinely distinct conceptualizations or alternative representations of similar themes.
5.4 Human Validation of Topic Quality
Our human evaluation by three independent domain experts provided essential confirmation that our computationally optimized topics align with substantive interpretations. The validation yielded robust inter-annotator agreement across multiple metrics. The Intraclass Correlation Coefficient (ICC(2,1) = 0.79) indicates very good reliability [koo_guideline_2016], demonstrating consistent evaluation across raters. Mean pairwise weighted Cohen’s kappa of 0.578 (moderate agreement) accounts for the ordinal nature of ratings while showing substantial consensus. Most strikingly, annotators achieved perfect agreement within a one-point tolerance (100%), indicating high consistency in perceived topic quality even when exact ratings differed marginally.
Six of nine topics received high ratings (mean > 4.0), with two achieving perfect scores (5.0). Only one topic received a lower rating (mean = 2.0), signaling potential for refinement. Topic size showed no correlation with coherence ($\rho$ = -0.213, p = 0.583), confirming that ratings reflected semantic quality rather than prevalence.
These findings validate that our systematic computational pipeline produces topics that domain experts recognize as interpretable and meaningful, addressing longstanding concerns [nanda2023, steele2025] that topic modeling algorithms may generate statistically optimal but substantively incoherent clusters. The strong agreement across three independent raters further suggests that the identified topics capture genuine thematic structures rather than analyst-specific interpretations. However, our validation focused exclusively on the final 9 merged topics; the 49 fine-grained initial topics were not independently evaluated. Future work could examine whether human raters prefer fine-grained versus consolidated topic structures, and whether stability-optimized configurations yield more interpretable topics than coherence-optimized alternatives.
5.5 Limitations
Several limitations merit consideration. First, the substantial sensitivity of results to hyperparameter choices raises questions about generalizability. Our finding that min_cluster_size alone shifted stability from 0.376 to 0.700, coherence from 0.385 to 0.589, and topic count from 2 to 52 demonstrates that configuration choices profoundly affect outcomes. While our systematic exploration represents a more rigorous approach than typical practice, specific optimal configurations for Tunisian HPV vaccine focus groups may not transfer to focus groups on different topics, with different conversational dynamics, or in different cultural contexts.
@@ 5.5 Limitations
Second, our study is limited to French-language transcripts from Tunisia. The multilingual embedding model we employed may perform differently across languages, and the identified topic structures likely reflect culturally specific patterns of health discourse. Validation in other linguistic and cultural contexts is needed to assess the broader applicability of our methodological framework.
Third, our LDA baseline was tuned only on the number of topics following current practice [hagg2022], potentially underestimating LDA’s optimal performance. More extensive tuning of prior distributions (alpha, beta) and learning parameters could improve LDA coherence. More critically, we did not conduct parallel human evaluation of LDA topics, which would have provided direct empirical comparison of interpretability beyond algorithmic metrics.
Finally, we compared BERTopic only with LDA, not with other neural topic modeling approaches such as Top2Vec or CTM (Contextualized Topic Models). Egger and Yu [egger2022] found notable performance differences across methods depending on corpus characteristics, suggesting value in comprehensive benchmarking studies.
@@ 5.6 Implications for Focus Group Research
Despite these limitations, our work demonstrates the feasibility and value of applying rigorous topic modeling to focus group transcripts. To our knowledge, this is the first study to develop and validate a systematic topic modeling framework specifically for focus group discussions. This contribution is meaningful because focus groups remain a cornerstone methodology in qualitative health research, yet the labor-intensive nature of traditional coding limits the scale of synthesis possible. Computational approaches do not replace human interpretation but can augment it by efficiently identifying thematic patterns across multiple sessions.
Our findings suggest practical recommendations for researchers considering topic modeling for focus group analysis. First, invest in systematic hyperparameter exploration rather than accepting default configurations. Performance varies dramatically, and optimal settings depend on corpus characteristics. Second, prioritize stability assessment through bootstrap resampling. Coherence scores alone provide incomplete quality information. Third, use multiple complementary metrics (stability, coherence, coverage, granularity) and be transparent about tradeoffs in model selection. Fourth, validate computational outputs through human review by domain experts. Finally, consider hierarchical approaches that extract fine-grained topics for stability then consolidate for interpretability. This two-stage strategy can navigate the stability-coherence tradeoff effectively.
@@ 6 Conclusion
This paper presents a systematic, reproducible computational framework for applying BERTopic to focus group transcripts, contributing validated methodological practices to the computational social science toolkit. Through rigorous exploration of 27 hyperparameter configurations with bootstrap stability assessment and formal human validation, we establish that transformer-based topic modeling can reliably extract interpretable themes from structured conversational data. The strong inter-rater agreement (ICC = 0.79) confirms that computational approaches, when properly validated, produce outputs aligned with expert domain understanding.
Our work addresses a critical gap: while topic modeling has been extensively applied to social media and online text, focus group discussions, a cornerstone of qualitative research, have received limited computational attention. By demonstrating feasibility on transcripts with naturally imbalanced topic distributions and relatively small corpus size (1,076 utterances), we extend neural topic modeling to structured research settings where traditional coding remains dominant. The transparent documentation of all modeling choices, evaluation procedures, and quality tradeoffs provides a reproducible template that researchers can adapt to their own focus group contexts.
This methodological foundation enables substantive analysis of HPV vaccine perceptions in Tunisia, where our identified themes reveal specific patterns in health discourse, including vaccine hesitancy drivers, information source credibility, gender-specific concerns, and trust in health institutions. These findings will inform targeted communication strategies and policy interventions in the forthcoming companion paper. More broadly, our framework supports the integration of computational efficiency with interpretive depth, allowing researchers to scale qualitative synthesis across multiple sessions while maintaining the contextual understanding that characterizes rigorous social research. As focus groups remain central to health research, market research, and social science inquiry, validated computational methods that respect conversational data characteristics hold significant potential for advancing both methodological practice and substantive understanding.
@@ Appendix A Appendix: Representative Topic Examples
Table 4 provides representative utterances for each of the 9 merged topics identified by our hierarchical BERTopic model, presented in both original French and English translation. These examples illustrate the thematic content and semantic coherence of each topic cluster.
@@
0 - COVID-19 vaccine side effects and trust erosion Pardon, mais d’après mon expérience personnelle et ce que j’entends, la majorité des gens ont eu des troubles de mémoire après le vaccin contre le Covid. Je suis devenu quelqu’un qui oublie beaucoup, et d’autres confirment aussi ce constat. Moi, sincèrement, avant la crise du COVID, j’avais une confiance totale dans notre système de santé et je n’ai jamais pensé que c’était nécessaire de poser des questions, surtout que la Tunisie a une longue tradition en matière de vaccination. Mais après le COVID, je n’ai plus confiance. Le vaccin contre la grippe, je l’ai fait, c’est un peu cher, cette année je ne l’ai pas fait. Sorry, but based on my personal experience and what I hear, the majority of people have had memory problems after the Covid vaccine. I have become someone who forgets a lot, and others also confirm this observation. Honestly, before the COVID crisis, I had complete trust in our healthcare system and I never thought it was necessary to ask questions, especially since Tunisia has a long tradition of vaccination. But after COVID, I no longer have trust. I got the flu vaccine; it’s a bit expensive. This year, I didn’t get it.
1 - Conservative norms and sexual health education Vous savez, la société tunisienne est conservatrice, même lorsque quelque chose comme ça se produit, on n’en parle pas, c’est honteux. Parce que les filles sont devenues plus libres, nous avons fait une campagne de sensibilisation dans les lycées pour expliquer aux filles et aux garçons l’importance de prendre un préservatif pour se protéger. Mais, nous avons été stigmatisés. Les gens ont dit que nous sommes en train d’encourager les filles à être encore plus libres. Alors que c’est juste de la prévention. Non, en 4ème, par exemple, on lui dit que le couple est marié et que la mère est enceinte. Mais cela n’explique pas l’acte sexuel. You know, Tunisian society is conservative; even when something like that happens, people don’t talk about it, it’s shameful. Because girls have become freer, we conducted an awareness campaign in high schools to explain to girls and boys the importance of taking a condom to protect themselves. But we were stigmatized. People said that we were encouraging girls to be even freer. Whereas it is just prevention. No, in 8th grade, for example, they are told that the couple is married and that the mother is pregnant. But that does not explain the sexual act.
2 - Historical trust in national vaccine programs Les vaccins sont considérés comme un programme réussi. Lorsque vous comparez la Tunisie à des pays d’Afrique où ils sont confrontés à de nombreuses maladies graves en raison du manque de vaccins nécessaires, nous devons apprécier ce que nous avons. La vaccination en général est une bonne chose car elle protège contre la poliomyélite, l’hépatite et ceux qui ne reçoivent pas les vaccinations nécessaires peuvent avoir des problèmes de santé plus tard. Dans le public, ils ont leur carnet de vaccination. Avant de rentrer à l’école, ils doivent être vaccinés. Vaccines are considered a successful program. When you compare Tunisia to African countries where they face many serious diseases due to the lack of necessary vaccines, we must appreciate what we have. Vaccination in general is a good thing because it protects against poliomyelitis, hepatitis, and those who do not receive the necessary vaccinations may have health problems later on. In public, they have their vaccination record. Before entering school, they must be vaccinated.
3 - Parental concerns about HPV vaccination Principalement pour protéger la santé de ma fille. Comme c’est la cadette, je voulais la protéger autant que possible. Quand tu lui dis que la maladie est liée au col de l’utérus et qu’elle ne concerne que les filles, certainement que ces données le feront hésiter et penser que le vaccin pourrait peut-être affecter la fertilité de sa fille. C’est pourquoi j’insiste sur l’importance de mener des campagnes de sensibilisation pour expliquer aux gens que les filles représentent la population la plus touchée par ce virus. Mainly to protect my daughter’s health. Since she is the youngest, I wanted to protect her as much as possible. When you tell him that the disease is related to the cervix and only affects girls, these facts will certainly make him hesitate and think that the vaccine might possibly affect his daughter’s fertility. That is why I emphasize the importance of conducting awareness campaigns to explain to people that girls are the population most affected by this virus.
4 - Information sources and communication channels Les médecins doivent être formés à communiquer efficacement avec les parents. La télévision atteint un plus grand nombre de citoyens et on ne peut pas mentir à la télévision. Les gens dans les réseaux sociaux numériques publient des informations fausses et les gens finissent par les croire. Par exemple, récemment, ils ont dit qu’il y avait de la neige à Ain Draham alors qu’il n’y avait ni neige ni rien. C’est de la désinformation. Doctors must be trained to communicate effectively with parents. Television reaches a larger number of citizens and one cannot lie on television. People on digital social networks publish false information and people end up believing it. For example, recently, they said there was snow in Ain Draham when there was neither snow nor anything. This is misinformation.
5 - Personal experiences with cancer and its social impact Je connais quelqu’un qui en a souffert et qui a tout enlevé, elle avait 38 ans, elle en a été affectée et déprimée. Elle est mariée et a un enfant, c’est ma cousine, elle a été hospitalisée, elle a subi une opération pour enlever son utérus. Il y a aussi une maladie très répandue qui est le cancer du foie. J’ai subi deux opérations chirurgicales pour l’enlever. N’y a-t-il pas de vaccination contre cela ? J’ai une voisine qui est tombée malade de cette maladie et son mari l’a divorcée. Les hommes ne soutiennent pas les femmes dans ces situations. I know someone who suffered from it and had everything removed; she was 38 years old, and it affected her deeply and made her depressed. She is married and has a child; she is my cousin. She was hospitalized and underwent surgery to have her uterus removed. There is also a very common disease which is liver cancer. I have undergone two surgical operations to remove it. Is there no vaccination against it? I have a neighbor who fell ill with this disease and her husband divorced her. Men do not support women in these situations.
6 - Prevention practices and barriers to care Moi, je ne veux pas faire le frottis car j’ai peur. J’ai eu une mauvaise expérience après mon accouchement dans un hôpital public et je ne veux pas que quiconque me touche dans ma zone intime. Oui, je consulte mon médecin chaque année pour un contrôle. Il y a des remèdes naturels que j’utilise quand mon enfant est malade, comme le miel et l’eucalyptus. I don’t want to have the smear test because I’m afraid. I had a bad experience after my delivery in a public hospital and I don’t want anyone to touch me in my intimate area. Yes, I see my doctor every year for a check-up. There are natural remedies that I use when my child is sick, such as honey and eucalyptus.
7 - Polarized attitudes towards official public health sources Je ne sais pas, je n’ai pas confiance dans l’ensemble du système. Nous faisons confiance aux vaccins approuvés par l’État et le ministère de la Santé. Moi, je fais confiance à la page officielle du ministère de la Santé. I don’t know, I don’t trust the whole system. We trust vaccines approved by the state and the Ministry of Health. I trust the official page of the Ministry of Health.
8 - General disease awareness and knowledge gaps C’est la première fois que j’entends parler de cette maladie. L’hygiène est l’une des principales causes de cette maladie. Je n’ai entendu parler d’aucun cas, je sais seulement que c’est une maladie maligne. This is the first time I have heard of this disease. Hygiene is one of the main causes of this disease. I have not heard of any cases; I only know that it is a malignant disease.
@@
Table 4: Representative utterances for each of the 9 merged topics with original French text and English translations.

### v2

@@ 5.3 Comparison with LDA Baseline
The substantial performance gap between BERTopic and LDA (18% higher coherence, 7 vs. 3 topics) likely reflects fundamental architectural differences. LDA’s bag-of-words assumptions treat documents as unordered token collections, relying on statistical co-occurrence patterns to infer latent topics. In contrast, BERTopic’s transformer-based sentence embeddings capture contextual semantics, preserving meaning that depends on word order and discourse context.
For conversational focus group data, this distinction proves particularly important. The same phrase may reference different concepts depending on surrounding discourse. Contextual embeddings can distinguish these semantic nuances, while bag-of-words representations collapse them into undifferentiated co-occurrence statistics. The LDA baseline’s optimal configuration at only 3 topics suggests that without contextualized representations, the model struggles to differentiate finer semantic distinctions, producing overly broad topics that sacrifice thematic resolution for coherence.
However, our comparison has limitations. Following standard practice [hagg2022], we tuned only the number of topics for LDA while maintaining default hyperparameters for Dirichlet priors and learning parameters. More extensive hyperparameter exploration could potentially improve LDA performance, though we consider our approach representative of typical applied usage. Additionally, we did not compare against other neural topic modeling approaches such as Top2Vec or CTM, limiting our ability to assess BERTopic’s performance relative to alternative contextualized methods.
5.4 Limitations and Future Directions
@@
Several limitations merit consideration. First, our study is restricted to French-language focus group transcripts from Tunisia. The multilingual embedding model we employed may perform differently across languages, and the identified topic structures likely reflect culturally specific patterns of health discourse. Validation in other linguistic and cultural contexts is needed to assess the broader applicability of our methodological framework.
Second, the substantial sensitivity of results to hyperparameter choices raises questions about generalizability. While our systematic exploration represents a more rigorous approach than accepting default settings, specific optimal configurations may not transfer to focus groups with different conversational dynamics or substantive domains.
Third, one large topic received poor coherence ratings from evaluators, indicating the model produced a heterogeneous catch-all category rather than a unified theme. This suggests limitations in BERTopic’s density-based clustering for highly diverse discourse, where some utterances may not fit coherent thematic patterns. Finally, we compared only against LDA rather than other neural methods, and our LDA baseline was tuned only on number of topics following standard practice. More comprehensive benchmarking across methods and fuller hyperparameter exploration would strengthen comparative conclusions.
Future work should examine whether the stability-coherence tradeoff generalizes across languages and data types, explore alternative stability metrics that better align with interpretability, and conduct systematic comparisons with other neural topic modeling approaches. Integration with manual qualitative coding frameworks could also clarify how computational topics relate to human-identified themes.
@@ 5.5 Implications for Focus Group Research
This work demonstrates the feasibility and value of applying systematic topic modeling to focus group transcripts. Focus groups remain a cornerstone methodology in qualitative health research, yet the labor-intensive nature of traditional manual coding limits the scale of synthesis possible across multiple sessions. Computational approaches do not replace human interpretation but can augment it by efficiently identifying thematic patterns, particularly valuable for large-scale focus group studies where manual analysis of hundreds of transcripts becomes prohibitive.
Our findings suggest several practical recommendations for researchers considering topic modeling for focus group analysis. First, invest in systematic hyperparameter exploration rather than accepting default configurations, as performance varies dramatically depending on corpus characteristics. Second, evaluate models across multiple complementary metrics (stability, coherence, granularity, and coverage) and transparently document the tradeoffs inherent in final selection decisions. Third, validate computational outputs through independent human review by domain experts using structured protocols; we recommend at least three raters evaluating representative documents per topic on standardized coherence scales. Fourth, apply minimal preprocessing for transformer-based models, preserving stop words and contextual elements that bag-of-words approaches would discard. These practices balance computational efficiency with the interpretive rigor qualitative research demands, providing a principled framework for augmenting traditional focus group analysis with algorithmic support.
@@ 6 Conclusion
This paper presents a systematic, reproducible computational framework for applying BERTopic to focus group transcripts, contributing validated methodological practices to the qualitative research toolkit. Through rigorous exploration of 27 hyperparameter configurations with bootstrap stability assessment and formal human validation, we establish that transformer-based topic modeling can reliably extract interpretable themes from structured conversational data. The strong inter-rater agreement (ICC = 0.700) confirms that computational approaches, when properly validated, produce outputs aligned with expert domain understanding.
Our work addresses a critical gap: while topic modeling has been extensively applied to social media and online text, focus group discussions, a cornerstone of qualitative research, have received limited computational attention. By demonstrating feasibility on transcripts with naturally imbalanced topic distributions and relatively small corpus size (1,075 utterances), we extend neural topic modeling to structured research settings where traditional coding remains dominant. The transparent documentation of all modeling choices, evaluation procedures, and quality tradeoffs provides a reproducible template that researchers can adapt to their own focus group contexts.
This methodological foundation enables substantive analysis of vaccine perceptions in Tunisia, where our identified themes reveal specific patterns in health discourse, including vaccine hesitancy drivers, information source credibility, and trust in health institutions. These findings will inform targeted communication strategies and policy interventions in the forthcoming companion paper. More broadly, our framework supports the integration of computational efficiency with interpretive depth, allowing researchers to scale qualitative synthesis across multiple sessions while maintaining the contextual understanding that characterizes rigorous social research. As focus groups remain central to health research, market research, and social science inquiry, validated computational methods that respect conversational data characteristics hold significant potential for advancing both methodological practice and substantive understanding.
@@ Appendix A Appendix: Representative Topic Examples
Table 5 provides representative utterances for each of the 7 topics identified by our BERTopic model, presented in both original French and English translation. These examples illustrate the thematic content and semantic coherence of each topic cluster.
@@
0 - General disease knowledge, health fears, and taboos Je pense que c’est un sujet délicat et on ne peut pas en parler avec tout le monde. Oui, exactement. Ma collègue m’a confié que ce qu’elle vivait dépassait ma compréhension. Par exemple, les douleurs qu’elle ressentait aux toilettes, les symptômes étranges… C’est une maladie existante mais rare. I think this is a sensitive topic, and you cannot talk about it with everyone. Yes, exactly. My colleague confided that what she was going through was beyond my understanding. For example, the pain she felt when using the bathroom, the strange symptoms… It is a real but rare disease.
1 - General vaccination importance Moi, je suis convaincue qu’il faut faire tous les vaccins, bien que je n’aie pas encore d’enfants. La question de la vaccination obligatoire est importante depuis toujours, car les établissements scolaires et d’autres institutions exigent des élèves, étudiants et employés ce document, le certificat de vaccination. La majorité d’entre nous suivent le carnet de vaccination. I am convinced that one should take all vaccines, even though I do not yet have children. The issue of mandatory vaccination has always been important because schools and other institutions require students and employees to provide this document: the vaccination certificate. Most of us follow the vaccination schedule.
2 - Trusted sources of health information Je prends l’information de tout médecin ou spécialiste, qu’il travaille dans le secteur public ou privé. Il ne faut pas dépendre des médecins. Chacun peut s’informer par ses propres moyens. Facebook est une source importante pour moi et j’ai confiance dans les informations publiées sur Facebook. Je ne suis pas d’accord avec elle. Le ministère de la Santé n’est pas n’importe qui. Il faut suivre les instructions du ministère. I take information from any doctor or specialist, whether they work in the public or the private sector. You should not depend on doctors. Everyone can inform themselves on their own. Facebook is an important source for me, and I trust the information published there. I do not agree with her. The Ministry of Health is not just anyone. You must follow the ministry’s instructions.
3 - Personal experiences with vaccination Moi, je ne l’ai pas reçu car j’étais enceinte à l’époque de la vaccination et j’ai attrapé le virus. J’ai beaucoup souffert et j’ai eu de gros problèmes respiratoires. Je n’étais pas convaincue, mais j’ai dû me faire vacciner. Le vaccin contre la grippe, je l’ai fait, c’est un peu cher, cette année je ne l’ai pas fait. I did not receive it because I was pregnant at the time of vaccination, and I caught the virus. I suffered a lot and had severe respiratory problems. I was not convinced, but I had to get vaccinated. I got the flu vaccine; it is a bit expensive. This year, I did not get it.
4 - Private vs Public sector debate C’est injuste. Vous pensez que c’est normal que je paie cette somme pour un vaccin que j’ai normalement le droit d’obtenir gratuitement ! C’est mon droit ! Mais il y a plusieurs lacunes dans le secteur public qu’il faut reconnaître. Le secteur privé dans le domaine de la santé est devenu supérieur au secteur public… Parce que ma fille a suivi sa grossesse chez un médecin privé et a accouché dans une clinique privée. Elle n’a donc pas eu recours au secteur public dès le début, choisissant de surveiller son enfant chez un pédiatre privé. It is unfair. Do you think it is normal for me to pay this amount for a vaccine that I have the right to obtain for free? It is my right! But several shortcomings in the public sector must be acknowledged. The private health sector has become superior to the public sector… Because my daughter followed her pregnancy with a private doctor and gave birth in a private clinic. She therefore did not use the public sector from the beginning, choosing to monitor her child with a private pediatrician.
5 - COVID-19 vaccine adverse effects Le Covid a vraiment changé la donne. Personnellement, après avoir reçu le vaccin contre le Covid, mon immunité a diminué. Avant, je guérissais d’une grippe en une semaine, mais maintenant, cela prend au moins un mois. Moi, mon ophtalmologue m’a critiqué car j’ai fait le vaccin contre le Covid. J’ai eu des problèmes dans mes yeux après avoir fait ce vaccin. Le vaccin contre le COVID provoque Alzheimer. Covid really changed things. Personally, after receiving the Covid vaccine, my immunity decreased. Before, I recovered from the flu in one week, but now it takes at least a month. My ophthalmologist criticized me for taking the Covid vaccine. I had eye problems after receiving it. The Covid vaccine causes Alzheimer’s.
6 - Erosion of institutional trust post pandemic Après la pandémie de COVID, l’opinion publique est devenue plus prudente. Non, auparavant, nous avions une confiance aveugle dans le ministère de la Santé, mais les failles sont apparues après la pandémie de COVID, et notre confiance dans le ministère de la Santé a diminué… Le COVID a créé ce problème, avant il n’y avait pas de précautions contre ceci ou cela. Depuis que nous avons appris les risques du vaccin et qu’il peut provoquer une crise cardiaque, les gens ont peur. After the COVID pandemic, public opinion became more cautious. No, previously we had blind trust in the Ministry of Health, but flaws became apparent after the COVID pandemic, and our trust in the Ministry of Health decreased… COVID created this problem; before, there were no precautions about this or that. Since we learned about the risks of the vaccine and that it can cause a heart attack, people are afraid.
@@
Table 5: Representative utterances for topics with original French text and English translations.